OpenStudy (pphalke):

PLEASE HELP!! nico is saving money for his college education. he invests some money at 6% and $1900 less than the amount at 5%. the investments produced a total of $191 interest in 1 yr. how much did he invest at each rate?

OpenStudy (pphalke):

PLEASE HELP ME

OpenStudy (anonymous):

Is the interest simple of compounding? That'll be key to calculating this one. If it is compound, what is the compounding frequency?

OpenStudy (pphalke):

thats whats confusing me. i dont know how to do this

OpenStudy (anonymous):

Let's look at the simple interest case. It is more approachable.

OpenStudy (pphalke):

okay

OpenStudy (pphalke):

its asking for the interest at each rate 5% and 6%

OpenStudy (pphalke):

i gave up on this problem so i didnt know what to do @eashmore

OpenStudy (anonymous):

Simple interest always pays interest against the original amount invested. Let's say I invested $100 with a bank with a simple interest rate of 5% a year. At the end of the first year, I would have $105 in the bank. That's because 5% of 100 is 5, so they pay me $5. The formula to calculate interest paid on a load (or paid to a savings account) is\[I = P * R * T\] I is the interest paid. In my example about, that is $5. P is the principle (or original amount borrowed or saved). In my example is $100. R is the interest rate (expressed as a decimal). In my example is 5%. T is the time. We want to use the same unit of time as the interest is paid. In my example, we get paid interest every year. The question gives us the Interest (I) earned. We can assume interest is paid our annually so Time (T) will be 1. It also gives us some idea about how much was invested at different interest rates. If x is the amount we invested, we will have two equations. I_1 is amount of interest earned at 6% and I_2 is the amount of interest earned at 5%. \[I_1 = x * 0.06 * 1\]\[I_2 = (x-1900)*0.05*1\] Together, these add to $191. So we get\[I_1 + I_2 = 191\]We need to substitute in our equations\[x * 0.06 * 1 - (x-1900) * 0.05 * 1 = 191\]

OpenStudy (anonymous):

My last equation has a typo. It should be a plus\[x*0.06*1 + (x-1900) * 0.05 *1 = 191\]

OpenStudy (anonymous):

Now we can combine like terms. \[0.06x + 0.05x - (1900*0.05) = 191\]\[0.11x - (1900*0.05) = 191\]

OpenStudy (anonymous):

If we solve for x we determine how much was invested at 6%. The amount invested at 5% will be x - 1900.

OpenStudy (pphalke):

i dont know how to solve for x to get the amount invested at 6% im stumped there

OpenStudy (pphalke):

@eashmore

OpenStudy (anonymous):

Take 1900*.05. What do you get?

OpenStudy (pphalke):

950

OpenStudy (pphalke):

is that for 6% or 5%

OpenStudy (anonymous):

We have to solve the equation. \[0.11x - (1900*0.05) = 191\]

OpenStudy (anonymous):

THe first step is to find the product within the parenthesis. This is 1900 times 0.05.

OpenStudy (anonymous):

You should get 95.

OpenStudy (anonymous):

Now, we have \[0.11x - 95 = 191\]

OpenStudy (anonymous):

We need to add 95 to both sides.

OpenStudy (anonymous):

Now we have \[0.11x = 286\]

OpenStudy (anonymous):

Divide by 0.11\[x = 2600\]

OpenStudy (anonymous):

X is amount invested at 6%. In my earlier expression, we defined what was invested at 5% to relate to what was invested at 6% as (x-1900)

OpenStudy (pphalke):

is the 2600 for 5% or 6%

OpenStudy (anonymous):

This helps us get down to one variable to solve for.

OpenStudy (pphalke):

so the 2600 is for six percent???

OpenStudy (anonymous):

The problem states that some amount is invested at 6%. We will call this x. The problem states that he invests 1900 less than this amount at 5%. Therefore, the amount invested at 5% is equal to x-1900. We are relating the two investment amounts as given in the problem statement.

OpenStudy (anonymous):

We solved the equation and determined that x was 2600. Which is the amount invested at 6%. What is the amount invested at 5% if it is 1900 less than the amount invested at 6%?

OpenStudy (pphalke):

subsitute x for 2600 and then subtract 2600-1900???

OpenStudy (pphalke):

if so then 700 is what i get but plz let me know if thats right

OpenStudy (pphalke):

@eashmore am i right if 2600-1900 is what i have to do then i get 700 is that right that he invested 700 at 5%??

OpenStudy (anonymous):

Correct.